# Click Without Compromise: Online Advertising Measurement via Per User Differential Privacy

### v2

@@ 2 Ad Measurement and Differential Privacy
$\mathbf{d}$: The difference between two neighboring dataset, $\mathbf{d}={{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}-\tilde{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}}$.
@@
$\mathbf{Q}$: Query workload matrix, it’s a stack of query vectors.
@@
$\mathbf{z}$: Random variable drawn from a Gaussian distribution.
@@
$l$: Number of days using differentially private quantile.
$p$: The quantile percentage used in the private quantile.
$s^{\uparrow},s^{\downarrow}$: Scaling parameters for updating the user contribution bound.
@@ 4.4 Estimate User Contribution Bound
12:  Output: The selected user contribution bound based on $p$ quantile.
23:  $X$ = [].
4:  for Each user $u_{i}$ in $D$ do
5:     $c_{i}\leftarrow$ the number of records from $u_{i}$.
6:     $X$.append($c_{i}$).
37:  end for
8:  Sort $X$ in ascending order. $X=[c_{1},c_{2},\cdots,c_{k}]$.
49:  Replace $c_{i}>\Lambda$ with $\Lambda$. Define $c_{0}=0$ and $c_{k+1}$.
10:  for  $i=1$ to $k$  do
11:     Set $y_{i}=(c_{i+1}-c_{i})\exp(-\epsilon|i-p*k|/2)$.
12:  end for
14:  $c\leftarrow$ uniformly sampled from $[c_{i},c_{i+1}]$.
15:  Return $c$.
Algorithm 2 PrivateQuantile
@@
Algorithm PrivateQuantile 2 satisfies $\rho_{2}$-zCDP with privacy paramter $\rho_{2}=\min(\frac{1}{8}\epsilon^{2},\frac{e^{\epsilon}-1}{e^{\epsilon}+1}\epsilon)$.
@@ 4.5 Track the Change of Contribution Bound
Algorithm UpdateBoundSVT 4 satisfies $\rho_{3}$-zCDP with $\rho_{3}=\frac{e^{\epsilon}-1}{e^{\epsilon}+1}\epsilon$.
@@ 5 Discussion
In AdsBPC, we add independent noise to the data vector with different noise scales. A natural question arises: why not use the matrix mechanism to optimize these noise scales? Our reasons are three-fold. Complexity: Solving the exact matrix mechanism under user-level differential privacy is NP-hard [14]. We compare AdsBPC with two variants of the matrix mechanism that approximate the user-level constraint in our experiments. AdsBPC consistently performs better, while also being simpler—eliminating the need for matrix optimization. Optimality for Small $n$: When the number of days $n$ is small and the objective function is the root mean squared error, we calculated the optimal matrix factorization using the code in [14]. Surprisingly, the resulting strategy matrix is diagonal, yielding the same objective value as AdsBPC. See the Appendix 8.8 for detailed numerical results. Whether AdsBPC achieves optimal noise scales in general remains a question for future research. Flexibility: The matrix mechanism is designed to minimize the sum of variances, limiting it to a single objective function. AdsBPC, on the other hand, offers greater flexibility to tailor to various objectives.
@@ 6.3 Datasets
TABLE X: The query is the sliding window sum, the objective (error metric) is the max variance (divied by $10^{4}$). The total privacy budget is $\rho=1$. The number of days is $n=31$.
@@ 6.4 Experiments on Different Queries and Objective Functions
In AdsBPC and MMBPC, we use the differentially private quantile algorithm for the first $l=7$ days, the quantile percentage is $p=99\%$, the bounding factor is $\Lambda=32$. In the SVT algorithm, the scale parameters are ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s}}^{\uparrow}=1.3$ and ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s}}^{\downarrow}=0.8$, the thresholds are $T^{\uparrow}=T^{\downarrow}=50$, the maximum number of reports is $k=7$. For the allocation of privacy budget, we set $\rho_{1}=0.7\rho$, $\rho_{2}=0.15\rho/l$ and $\rho_{3}=0.15\rho$.
@@ 6.5 Experiments on Different Parameters
Experiments on Different Ad Campaign Durations. In this section, we study the performance of different algorithms with respect to the number of days $n$, $n$ is chosen from [7, 15, 31, 63, 127]. The dataset we use is the Criteo dataset with LTA. The total privacy budget is set as $\rho=1$. Figure 2 illustrates the performance of various algorithms. AdsBPC outperforms others consistently.

### v3

@@ 2 Ad Measurement and Differential Privacy
$\mathbf{d}$: The difference between two neighboring dataset.
@@
$\mathbf{Q}$: Query workload matrix.
@@
$\mathbf{z}$: Random variable.
@@
$l$: Number of days using DP-quantile.
$p$: The quantile percentage used in the DP-quantile.
$s^{\uparrow},s^{\downarrow}$: Scaling parameters.
@@ 4.4 Estimate User Contribution Bound
Algorithm PrivateQuantile 4 satisfies $\rho_{2}$-zCDP with privacy paramter $\rho_{2}=\min(\frac{1}{8}\epsilon^{2},\frac{e^{\epsilon}-1}{e^{\epsilon}+1}\epsilon)$.
@@ 4.5 Track the Change of Contribution Bound
Algorithm UpdateBoundSVT 3 satisfies $\rho_{3}$-zCDP with $\rho_{3}=\frac{e^{\epsilon}-1}{e^{\epsilon}+1}\epsilon$.
@@ 5 Discussion
In AdsBPC, we add independent noise to the data vector with different noise scales. A natural question arises: why not use the matrix mechanism to optimize these noise scales? Our reasons are three-fold. Complexity: Solving the exact matrix mechanism under user-level differential privacy is NP-hard [15]. We compare AdsBPC with two variants of the matrix mechanism that approximate the user-level constraint in our experiments. AdsBPC consistently performs better, while also being simpler—eliminating the need for matrix optimization. Optimality for Small $n$: When the number of days $n$ is small and the objective function is the root mean squared error, we calculated the optimal matrix factorization using the code in [15]. Surprisingly, the resulting strategy matrix is diagonal, yielding the same objective value as AdsBPC. See the Appendix B for detailed numerical results. Whether AdsBPC achieves optimal noise scales in general remains a question for future research. Flexibility: The matrix mechanism is designed to minimize the sum of variances, limiting it to a single objective function. AdsBPC, on the other hand, offers greater flexibility to tailor to various objectives.
Next, we discuss the differences in DP-quantile usage between DP-SGD [3] and advertising systems. While DP-quantile has been applied to ML training [5], the advertising scenario presents distinct challenges. In ML training, DP-quantile primarily determines the gradient norm, and [5] shows that even a minimal privacy budget can yield benefits. However, in advertising, conversion measurement accuracy is highly sensitive to the user contribution bound. A low privacy budget for estimating this bound introduces significant noise. If the bound is too small, excessive data cutoff causes high bias; if too large, increased noise leads to higher variance. Thus, allocating a sufficient privacy budget for DP-quantile is crucial. However, daily application depletes the budget. To mitigate this, we exploit the stability of user contribution quantiles, updating them only when significant changes occur. Our novel SVT-based approach tracks users exceeding a threshold, reducing budget consumption compared to daily DP-quantile updates.
Lastly, we distinguish prior work on DP-SGD using matrix mechanisms [14, 15, 17], emphasizing the fundamental differences between ML training and ads measurement, particularly in how user-level privacy and neighboring datasets are defined. In [14], a neighboring dataset is defined such that two samples from the same user must be separated by at least b steps. This constraint is valid in ML training, where sample order can be controlled. However, in ads measurement, this constraint does not hold, as users may have consecutive conversions, and the order of their data cannot be controlled. As noted in [15], solving a user-level DP problem without specific ordering constraints—such as in our case—is NP-hard. Both [14] and [15] introduce ordering constraints to make the matrix mechanism feasible. It is also important to note that [17] does not implement a user-level DP mechanism. The key intuition behind our approach comes from the unique characteristics of ads measurement. Unlike ML training, where updates depend on previous iterations, user contributions on different days can be independent—i.e., for a random user, the probability of contributing each day is equal. Furthermore, the sensitivity in our setting differs from that in ML training due to the distinct definitions of neighboring datasets. In our setting, a neighboring dataset is defined by replacing a user’s conversions, with a maximum of r conversions allowed per day. This difference in sensitivity makes our approach specifically tailored to ads measurement.
@@ 6.3 Datasets
TABLE X: The query is the sliding window sum, the objective (error metric) is the max variance (divided by $10^{4}$). The total privacy budget is $\rho=1$. The number of days is $n=31$.
@@ 6.4 Experiments on Different Queries and Objective Functions
In AdsBPC and MMBPC, we use the differentially private quantile algorithm for the first $l=7$ days, the quantile percentage is $p=99\%$. In the SVT algorithm, the scale parameters are ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s}}^{\uparrow}=1.3$ and ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s}}^{\downarrow}=0.8$, the thresholds are $T^{\uparrow}=T^{\downarrow}=50$, the maximum number of reports is $k=7$. For the allocation of privacy budget, we set $\rho_{1}=0.7\rho$, $\rho_{2}=0.15\rho/l$ and $\rho_{3}=0.15\rho$. The experiment for time complexity is in Appendix A.1. The experiment for parameter tuning is in Appendix A.3. The ablation study is in Appendix A.4
@@ 6.5 Experiments on Different Parameters
Experiments on Different Ad Campaign Durations. In this section, we study the performance of different algorithms with respect to the number of days $n$, $n$ is chosen from [7, 15, 31, 63, 127]. The dataset we use is the Criteo dataset with LTA. The total privacy budget is set as $\rho=1$. Figure 2 illustrates the performance of various algorithms. AdsBPC outperforms others consistently. The experiment for $n=365$ is in Appendix A.2.
7 Limitations and Future Work
Our approach currently assumes the presence of a trusted third party, similar to existing multi-touch attribution (MTA) systems where entities like advertising service providers (e.g., Criteo) or mobile measurement partners fulfill this role. However, reliance on a centralized third party introduces privacy risks, as some have faced regulatory scrutiny for mishandling user data. While our proposed algorithm with differential privacy (DP) mitigates inference attacks and cross-tracking risks, it does not fully eliminate concerns if a third party acts maliciously. As future work, we plan to explore decentralized alternatives to reduce or eliminate this dependency. Techniques such as Trusted Execution Environments (TEE) can shift trust from third parties to secure hardware providers, ensuring that only approved computations are executed. Alternatively, Multi-Party Computation (MPC) offers mathematically proven security without requiring a trusted intermediary. However, practical challenges remain in designing efficient protocols and ensuring DP guarantees within an MPC framework. Addressing these challenges will be a key direction for future research.